Given cscθ = 3, find cos(θ).
1
Expert's answer
2013-01-30T09:32:23-0500
3=csc x =1/sin x
sin x = 1/3
cos^2( x ) = 1- 1/9 = 8 / 9

cos x= 2*sqrt(2) /3
or
cos x= -2*sqrt(2) /3